# Audio-Language Models for Audio-Centric Tasks: A survey

### v1

@@ VI-C Benchmarks
VII Challenges and Future Directions
In this section, we will analyze some challenges currently faced by ALM research and humbly share possible research directions.
From a data perspective, there are three challenges and potential research directions.
1) Audio Sources: The current audio-language datasets have limited sources of audio samples [18], mainly including AudioSet [161], FreeSound [162], and BBC Sound Effects [163]. When using multiple audio-language datasets for model training, it is often limited due to the overlap of data sources, restricting the scaling up of model performance. In addition, the limited training data presents distribution shift issues [164], resulting in ALM lacking robustness and generalization ability when facing a variety of audio types and qualities. Compared with the data sources of LLM and VLM, the diversity of audio data is still limited [165]. We hope to develop methods for constructing online audio datasets to expand the nearly limitless scale of internet resources.
2) Text Sources: Currently, audio-language datasets focus on generating audio captions from known labeled datasets or manual annotation, hindering the acquisition of larger-scale datasets. The evolution of LLMs encourages more researchers to create ALM-suitable datasets via model generation methods. Despite their powerful performance, generated text may still contain inaccuracies like hallucinations [166]. The quality issue of generated datasets significantly restricts ALM development. Constructing dependable model generation methods is a crucial research direction needing further advancement. Additionally, the vast but noisy audio-related text on the internet, such as audio introductions and comments, remains overlooked.
3) Evaluation Benchmarks and Analysis: A unified model evaluation framework is essential for identifying model limitations and advancing the field. While there has been significant progress in ALM research, empirical analysis across a wide range of audio types, tasks, and models remains scarce. This gap is partly due to the limited availability of audio resources, as data used in evaluations should not overlap with the training set. Therefore, we encourage the construction of evaluation benchmarks with broader coverage beyond mainstream audio sources, from other public or private domains, and look forward to extensive empirical analyses based on these benchmarks.
From a model perspective, there are four challenges and potential research directions.
1) Unified multimodal encoder: Current ALMs mainly employ two separate networks for processing audio and language modal inputs. In the vision community, there is a proposition that unifying learning on the One Head architecture can enhance communication efficiency across data modalities, improve training efficiency, and alignment [167, 31]. However, there are relatively few related studies in the field of audio-language [35]. Compared to training methods, the model architecture is a more fundamental issue and deserves more effort to explore and improve.
2) Pre-trained Foundation Models: Literature has experimentally demonstrated that the choice of pre-trained audio and text encoders is crucial for generating robust representations and improving task performance [141, 22, 40]. However, with the rapid development of language models, the primary audio encoders are still pre-trained models such as PANNs [43] and AST [48]. We acknowledge that the open ALM community lacks a large, well-trained audio encoder to achieve performance improvements in ALM.
3) Continual Learning: Existing work highlights the use of larger-scale training data to enhance the pretraining performance of ALMs. However, as streamable data accumulates, pre-training from scratch requires increasingly large storage and computational resources, limiting the continuous growth of models. Moreover, when old data is unavailable, such pretraining from scratch is not permissible. We aspire for models to directly continue learning on new datasets without affecting their performance on old tasks and zero-shot predictions. This method of continual learning for ALMs has yet to be studied.
4) Efficient Learning: Compared to the natural language and vision communities, sources of audio are relatively limited, and pre-training ALMs on large-scale training data requires intensive computation, making sustainability a significant issue. We look forward to research on data-efficient ALMs that can train effective models with limited audio-text data. For example, learning not only from each audio-text pair but also utilizing more useful information from supervised learning between paired data [168]. Also employing active learning or curriculum learning methods are possible to make more efficient use of limited data [169]. Moreover, for audio tasks and transfer learning methods are orthogonal, we encourage combining ALMs with more transfer learning methods to explore a better trade-off between parameter-efficient learning and performance.
VIII Conclusion
ALMs can effectively utilize natural language to learn detailed descriptions of audio, achieving zero-shot predictions without any task-specific fine-tuning and demonstrating high performance on fine-tuned audio-centric tasks. Models combined with LLMs exhibit listening, thinking, and reasoning capabilities for complex audio and have achieved significant success in following audio instructions. This survey extensively reviews ALMs for audio tasks from multiple perspectives, including background, foundation, representation pre-training, downstream transfer, training and evaluation data. It abstracts the common characteristics of related work through illustration and formalization, and compares and summarizes mainstream research through tables. The survey covers a wide range of tasks in various audio types such as natural sounds, speech, and music, providing a clear panorama of the latest developments in ALMs. On this basis, we analyze the current challenges in this field and humbly propose potential research directions that we believe could benefit the future development of this emerging but very promising area.

### v2

@@ VI-C Benchmarks
VII Evaluations
This section details primary ALM evaluation methods, and results from common benchmarks to offer practical references.
VII-A Evaluation Methods
VII-A1 Zero-Shot Evaluation
This method evaluates ALMs’ open-set retrieval through similarity measurement between audio and text embeddings in a shared space. By treating labels as specialized language, it extends naturally to classification tasks on standard datasets like ESC50 [piczak2015esc50].
VII-A2 Linear Probe Evaluation
It evaluates audio representation quality by training a linear classifier (e.g., multilayer perceptron) on downstream tasks with frozen ALM features. While not optimal for specific tasks, this setup minimizes variables for fair representation comparison, commonly used in audio classification.
VII-A3 Supervised Fine-tune Evaluation
It examines model generalization by fine-tuning the entire model or its components on downstream tasks. The audio encoder, along with a task-specific head, is optimized end-to-end, and performance is validated on test sets with comparison to state-of-the-art (SOTA) models. It comprehensively assesses model adaptability by supporting fine-tuning across a wide range of tasks, such as classification, retrieval, and captioning [drossos2020clotho, kim2019audiocaps] and so on.
VII-A4 Instruction-following Evaluation
It measures the ability to understand and accurately follow human instructions, indicating task generalization in LALMs. This evaluation method can be considered a special type of zero-shot evaluation or supervised fine-tuning evaluation, depending on whether instruction tuning is performed.
VII-B Evaluation Results
Table VI summarizes zero-shot and linear probe results (mAP: mean Average Precision; Acc: Accuracy) on common audio classification tasks, widely used for evaluating contrastive learning representations and LALM benchmarks. Results confirm contrastive learning enables open-set capability, while adaptive training through a simple linear classifier significantly improves downstream task performance. Table VII shows supervised fine-tuning results on ATR, where R@k (Recall at rank k) reflects task adaptation capability. For instruction-following evaluation spanning diverse tasks and settings, readers are referred to established benchmarks [yang2024air, wang2025audiobench] for systematic LALM comparisons.
Notably, the field faces reproducibility and fairness challenges in model evaluation. As systematically revealed in [yang2025evaluationsurvey], YouTube-sourced datasets (e.g., AudioSet, AudioCaps) suffer from link rot, leading to inconsistent data compositions across studies and affecting result comparability. Additionally, data leakage is prevalent—unintended audio overlaps between different datasets (e.g., WavCaps and Clotho/AudioCaps) can inflate evaluation metrics if left unaddressed. Therefore, interpreting existing benchmark results requires accounting for these inconsistencies in evaluation setups. Future research should establish standardized data usage protocols and stricter de-duplication pipelines to ensure evaluation fairness and result reproducibility.
TABLE VI: Zero-Shot and Linear Probe Evaluation of ALMs on Audio Classification Tasks (%).
[…]
Zero-Shot
MS-CLAP [elizalde2023clap] 5.8 30.2 82.6 75.3 22.8 29.0 21.4 -
LAION-CLAP [wu2023large] - 45.9 91.0 77.0 23.1 47.2 35.3 30.0
MS-CLAP V2 [elizalde2024CLAP2] 10.2 48.5 93.9 82.3 30.0 58.4 58.1 -
WavCaps [mei2024wavcaps] 19.6 53.0 94.8 81.4 19.9 45.5 27.7 28.9
BLAT [xu2023blat] 10.5 31.3 80.6 77.3 - - - 14.9
Pengi [deshmukh2023pengi] 16.4 46.8 92.0 71.9 18.5 35.3 50.1 -
LTU [gong2024ltu] 18.7 46.3 83.1 - - 50.3 - -
M2D-CLAP [niizumi2024m2dclap] 27.2 40.8 75.5 72.4 17.7 75.2 23.4 -
Cacophony [zhu2024cacophony] - - 93.4 77.1 - - - 27.1
GLAP [dinkel2025glap] - 40.9 88.8 78.9 20.5 69.6 31.3 -
Linear Probe
MS-CLAP [elizalde2023clap] - - 93.8 84.2 54.4 79.3 68.2 -
LAION-CLAP [wu2023large] - - 97.3 86.9 54.6 84.3 72.2 -
MS-CLAP V2 [elizalde2024CLAP2] - - 97.7 88.4 62.5 82.3 80.5 -
WavCaps [mei2024wavcaps] - - 97.2 63.6 58.6 80.2 74.4 -
BLAT [xu2023blat] 38.7 32.4 95.8 85.7 - - - 42.9
Pengi [deshmukh2023pengi] - - 89.2 - 50.6 80.0 - -
M2D-CLAP [niizumi2024m2dclap] - - 96.3 88.8 73.4 84.1 78.0 -
[…]
  • Results are sourced from [niizumi2024m2dclap, zhu2024cacophony, xu2023blat, dinkel2025glap]; the best and second-best scores are in bold and underline, respectively.
TABLE VII: Supervised Fine-tune Evaluation of ALMs on Audio-Text Retrieval Tasks (%).
[…]
Metric R@1 R@5 R@10 R@1 R@5 R@10 R@1 R@5 R@10 R@1 R@5 R@10
MS-CLAP [elizalde2023clap] 33.5 70.4 80.2 47.8 80.2 90.7 16.2 39.6 51.4 23.6 46.7 60.3
LAION-CLAP [wu2023large] 35.1 71.8 83.9 45.8 80.9 91.6 18.2 42.5 54.4 25.7 51.5 63.4
MS-CLAP V2 [elizalde2024CLAP2] 35.6 - - 42.5 15.7 - - - - 22.9 - -
FLAP [yeh2023flap] 41.5 75.5 86.0 53.0 84.1 92.6 20.3 46.5 58.8 25.5 53.4 67.9
WavCaps [mei2024wavcaps] 39.7 74.5 86.1 51.7 82.3 90.6 19.5 45.2 58.2 23.4 50.9 63.4
Cacophony [zhu2024cacophony] 41.0 - 86.4 55.3 83.6 92.4 20.2 45.9 58.8 26.5 54.1 67.3
BLAT [xu2023blat] 38.2 - 85.1 47.5 - 87.6 13.7 - 48.9 17.9 - 50.9
AutoACD [sun2024auto] 39.5 - 85.4 53.7 - 91.7 15.3 - 52.1 17.7 - 52.6
T-CLAP [yuan2024tclap] 39.7 74.6 86.9 49.8 82.5 91.9 17.3 39.9 53.6 21.8 44.9 57.4
GLAP [dinkel2025glap] 41.7 - 86.1 54.4 - 91.1 19.4 - 58.3 21.8 - 61.5
[…]
  • Results are sourced from [niizumi2024m2dclap, zhu2024cacophony, xu2023blat, dinkel2025glap, yuan2024tclap]; the best and second-best scores are in bold and underline, respectively.
VIII Core Limitations and Concerns
Despite remarkable success in audio-centric tasks, deploying ALMs in real-world scenarios brings to light several critical limitations and concerns.
VIII-A Hallucination
Hallucination represents a fundamental limitation of LLMs that has also been identified in ALMs, where models generate content ungrounded in the source audio [kuan2024hallucination, arora2025landscape]. Beyond free-form captioning, models often affirm nonexistent sounds in structured QA, with ALMs showing higher hallucination rates than LLM cascades [kuan2025can]. Research attributes this limitation to an imbalanced capabilities: while models can generate accurate audio captions, they struggle to understand specific questions [kuan2024hallucination]. This disparity makes chain-of-thought prompting [wei2022chain] effective, as it leverages their captioning strength to support their weaker reasoning process for precise question answering.
VIII-B Vulnerabilities
While large-scale pretraining models demonstrate reasonable robustness to minor perturbations such as background noise or temporal stretching in general tasks [achiam2023gpt4], ALMs still exhibit significant security vulnerabilities against deliberately crafted adversarial attacks, particularly through jailbreak techniques that bypass their safety alignment. Early research efforts began with directly transferring text jailbreak methods via TTS systems [shen2024anything], followed by white-box attacks targeting open-source models [kang2025advwave]. More recent studies have explored spelled-letter prompts in TTS-generated audio [yang2025AchillesHeel], text-level sampling for prompt augmentation [hughes2024best], audio editing techniques manipulating tone and speed [xiao2025tune], and systematic benchmarking with audio characteristics and semantically-constrained adversarial perturbations [song2025ajailbreak]. Comprehensive evaluations of mainstream LALMs reveal that no single model demonstrates consistent robustness across all attack dimensions, highlighting critical security gaps (see comparative analysis in [song2025ajailbreak, hou2025evaluating]).
VIII-C Privacy Risks
ALMs introduce privacy risks by processing raw audio directly. Unlike cascaded pipelines, end-to-end models retain rich acoustic information (e.g., intonation, speaker identity), potentially enabling unauthorized user tracking via implicit voiceprint storage and inference of sensitive attributes like age and emotion [he2025model]. Environmental audio further risks exposing location, activities, or habits through soundscape monitoring, enabling covert large-scale surveillance beyond legal safeguards. Mitigation spans model-level techniques including safety alignment and in-context unlearning, and data-level protection such as anti-eavesdropping jamming [wang2025man].
VIII-D Bias
Web-scale training inherits and amplifies societal biases through imbalanced concept representation and spurious correlations [lin2024emobias, slaughter2023pre]. For instance, SER models may associate emotions with gender, while ASR performance degrades for underrepresented subgroups [koudounas2023subgroup]. Linguistic bias is also prevalent: the dominance of high-resource languages (e.g., English) in training data leads to significant performance disparities for low-resource languages, accents, and dialects [Peng2025SLLMsurvey, arora2025landscape]. Mitigations include data augmentation [zhang2023dataaug], focused data collection [koudounas2024prioritizing], and fairness-aware losses [zhang2022mitigating]. Most existing research operates on predefined subgroups such as race or valence, which has motivated exploration of alternative paradigms such as improving latent representations through contrastive learning [koudounas2024contrastivebias] and confidence-Oriented debiasing approach [tsai2025co].
VIII-E Training Costs
Scaling data and models creates significant accessibility and sustainability barriers. This is exemplified by contrastive pre-training, which requires large batches (e.g., FLAP used 72×64 GPUs for a 4608 batch size [yeh2023flap]). Although parameter-efficient methods like LoRA [hu2022lora] reduce LLM fine-tuning costs, LALMs still need extensive training to bridge modality gaps [gong2024ltu]. Representative examples include LTU (5.6M samples) and Audio Flamingo (5.9M samples; 8 A100s) [gong2024ltu, kong2024audioflamingo]. However, computational requirements for adapting ALMs vary significantly across downstream tasks. For example, TQ-SED [yin2024TQSED] fine-tunes only 0.25M parameters, demonstrating the potential for broader ALM deployment.
IX Future Research Directions
IX-A Towards Efficient and Scalable ALMs
Addressing the high computational cost of ALMs necessitates a shift towards more efficient and scalable learning methods. Promising directions encompass: Parameter-Efficient Architectures which leverage techniques such as model compression via distillation [liu2024audiolcm, paissan2024tinyclap] for compact edge models and adapter-based tuning for adaptation with minimal parameters; Data-Efficient Training utilizing active or curriculum learning [wan2023activesurvey] to maximize limited labeled data utility; and Lifelong Adaptation via continual learning [yang2024aqacontinual] that enable knowledge accumulation scalable to non-stationary data streams without catastrophic forgetting, thus avoiding expensive retraining cycles.
IX-B Strengthening Security Safeguards
As ALM generation capabilities advance, so do the risks of malicious use, particularly through deepfake audio creation. Future research must prioritize developing robust detection frameworks [li2023voiceguard] that can identify synthetic audio across diverse generation methods. Concurrently, applications such as voice assistants necessitate the integration of encryption, authentication, and access control technologies to provide technical safeguards against misuse [Simon2023AssistantSecurity]. Additionally, exploring adversarial training techniques and developing certified defenses against audio manipulation attacks will be crucial for building secure and trustworthy ALM systems.
IX-C Mitigating Algorithmic and Ethical Bias
Algorithmic and ethical bias in ALMs extends beyond simple demographic subgroups, manifesting in complex and systemic ways. For example, linguistic bias is particularly evident in voice assistants, where such performance disparities not only degrade the user experience for low-resource language users, such as those in developing regions, but also risk exacerbating the global digital divide [lima2019assistants]. Beyond language, acoustic biases related to voice timbre, age, or gender can similarly degrade performance for diverse user groups. Furthermore, socio-cultural biases may be embedded within training data, causing models to perpetuate stereotypes or fail to understand context dependent on local customs and values. Technical pathways to mitigate this spectrum of biases must be equally multifaceted. This includes exploring latent biases and enhancing fairness through techniques like contrastive learning adaptations [koudounas2024contrastivebias], and developing multilingual & multi-dialect training methods [babu2022xlsr] to learn more universal audio representations. Ultimately, these efforts are essential for deploying truly equitable ALMs in global domains like education, healthcare, and commerce.
IX-D Enhancing Real-World Application Readiness
Bridging the gap from laboratory to real-world deployment requires addressing domain-specific challenges alongside security and bias concerns. Voice assistants in practical deployment must ensure rapid response and privacy protection under noisy conditions. OpenAI’s Realtime API integrates ASR-LLM-TTS pipelines into a unified speech-to-speech architecture, significantly reducing latency [openai_realtime_api]. In healthcare, on-device deployment enables privacy-preserving services at the edge [shukla2025healthx]. ATR powers voice search in e-commerce, where systems must handle complex product names and code-mixing; Algolia’s solution integrates third-party speech transcription services (e.g., Google Cloud) to transcribe and optimize queries through data cleaning [algolia_voice_search]. In customer support scenarios, AQA systems must address the challenges of multi-turn dialogue understanding and intent recognition; Meituan’s WOWService employs multi-agent architectures for autonomous task management, significantly improving user satisfaction [cheng2025meituan]. Addressing these diverse and complex challenges is essential for the widespread adoption of next-generation audio-language models [gong2024ltu].
IX-E Developing Reliable and Comprehensive Evaluation Ecosystems
Building a reliable and comprehensive evaluation ecosystem for ALMs first requires addressing fundamental reproducibility issues and data leakage. As revealed in [wijngaard2024ald_survey], YouTube-sourced datasets suffer from link rot, leading to inconsistent data compositions; meanwhile, cross-dataset contamination is prevalent—for instance, substantial unintended audio overlaps exist between WavCaps and Clotho/AudioCaps, which can inflate metrics if left unaddressed. These issues severely compromise result comparability. Beyond these foundational concerns, ALM evaluation requires assessing capabilities beyond traditional metrics, including efficiency, safety, bias, and deployment constraints. While LLM research benefits from mature multi-dimensional benchmarks [zhao2023LLMsurvey], ALM work remains predominantly focused on cross-task performance. Although specialized evaluations for concerns like adversarial attacks are emerging [song2025ajailbreak], such benchmarks rarely simultaneously assess other helpfulness dimensions [yang2025evaluationsurvey]. Cultivating unified evaluation practices that encompass these critical dimensions is essential for guiding ALMs toward robust real-world deployment.
X Conclusions
ALMs leverage natural language to learn fine-grained audio representations, enabling zero-shot predictions and strong performance on fine-tuned tasks. By integrating LLMs, these models demonstrate advanced processing, understanding, and reasoning capabilities for complex audio. This survey provides the first systematic review of ALMs, integrating their development across speech, music, and natural sound within a unified audio-centric framework. We analyze the field from multiple perspectives: background, foundational architectures, pre-training, downstream transfer, data ecosystems, and evaluations. Through illustrations, formalizations, and comparative tables, we abstract common patterns in model architectures, training objectives, and evaluation methods, while mapping interdependencies between pre-training, transfer, datasets and benchmarks. Our review further identifies and discusses core limitations such as hallucination, security vulnerabilities, bias, and high training costs. Finally, we propose actionable directions to advance task performance and practical applications in complex audio scenes.